# ProST: Progressive Sub-task Training for Pareto-Optimal Multi-agent Systems Using Small Language Models

### v1

@@ Abstract
Biddut Sarker Bijoy*  Mohammad Saqib Hasan*  Pegah Alipoormolabashi* Avirup SilAruna Balasubramanian*Niranjan Balasubramanian* *Stony Brook University  IBM Research AI *{bbijoy, mdshasan, palipoormola, arunab, niranjan}@cs.stonybrook.edu,   avi@us.ibm.com
@@ 1 Introduction
Solving complex problems requires a wide-range of capabilities including planning sub-tasks, acting in an environment via coding, reasoning about the actions, recovering from errors, as well as tracking and managing the execution of sub-tasks. LLMs with their strong coding and reasoning abilities and advances in handling long contexts have shown impressive breakthroughs on such complex problems (e.g., AppWorld (Trivedi et al., 2024) and SWE-Bench (Jimenez et al., 2024)). However, these large models incur high computational and API costs. Another viable alternative is to use multi-agent solutions with smaller language models [^1] Small and large are subjective terms. For this paper we use SLM for models with fewer than 40B parameters. (SLMs). The core idea is to identify specialized roles in solving these complex tasks (e.g, orchestration, code writing, and critiquing) and use separate SLMs (or SLM calls) for each role. Belcak et al. (2025) argue that SLMs are much more cost-effective and practical for training, adapting, and deploying multiple specialized experts for different agentic routines.  Shen et al. (2024) show these multi-agent solutions with SLMs can be effective for complex problems requiring tool use via APIs.
We consider two challenges that arise in this context. First, for SLMs, even specialized roles can be difficult to learn. Complex problems often have multiple subtasks resulting in long trajectories. Despite their improving capabilities, SLMs still struggle to learn all subtasks under standard fine-tuning. Second, while the per-token compute depends on the size of the models, the overall computational cost also depends upon the nature of their solutions (e.g their length). Therefore, it is important to explicitly consider both the effectiveness and efficiency of these solutions to better understand the trade-offs.
@@
We evaluate these ideas using the Qwen family of models. We distill training trajectories for a single agent from a frontier model (Claude-3.7-Sonnet). Then, we translate these into role-specific multi-agent training trajectories for three agents, namely, an Orchestrator (planner), an Executor (coding) agent, and a Critic agent. Our evaluations yield multiple interesting observations AppWorld tasks: (i) While larger models tend to have higher effectiveness, multi-agent solutions can achieve better efficiency versus effectiveness trade-offs. (ii) Progressive Subtask Training yields a better Pareto-front compared to standard fine-tuning. (iii) Stronger (higher compute) agents for planning are more effective (and efficient) than using stronger coding and critic agents when finetuning using Progressive Subtask Training.
@@ 2 Related Work
SLMs are being used as agents due to lower computational cost, but struggle with complex planning and tool use (Shen et al., 2024), skills essential for agents. Hence, many strategies are proposed: separation of roles to only tackle small, concrete tasks (Shi et al., 2024; Qiao et al., 2024); distilling training data from powerful LLMs(Shi et al., 2024), Direct Policy Optimization (DPO)(Feng et al., 2025); online learning by generating data from unsuccessful attempts(Qi et al., 2025), masking loss from erroneous steps(Fu et al., 2025); learning only from critical steps (Chen et al., 2025d). Contrary to these works, we formulate ProST which optimizes SLMs, as part of multi-agent design, by initiating with important parts of the trajectory and iteratively increasing steps, a strategy never tried for agents.
@@
Many works propose techniques to make multi-agents more robust. These techniques include dynamically reducing redundant agents and messages (Wang et al., 2025b; Zhang et al., 2024), iterative answer refinement through feedback(Chen et al., 2025b), optimizing collaboration strategies(Wang et al., 2025a; Zeng et al., 2025) and communication protocols (Xiao et al., 2025), and test-time search for effective tool planning (Chen et al., 2025a). Recent works propose using cost-effective SLMs for multi-agent setups: Erdogan et al. (2025) uses smaller32-B models as planning and execution subagents; Belcak et al. (2025) argue for SLM multi-agents based on their improved capabilities, cost, ease of adaptability, and deployment. Our work introduces ProST, a training algorithm for SLMs in multi-agent setups that iteratively trains on subtasks. Unlike previous works, we show how ProST produce SLMs effective at solving tasks as well as computationally efficient.
@@
Recent works have sought to assess and improve the computational efficiency of multi-agent systems. For example, Wang et al. (2025b) use token consumption as a metric to prune redundancies in multi-agent collaborations. Yue et al. (2025) show, via Pareto fronts, that their language model routing strategy in multi-agent systems reduces dollar cost while improving accuracy. Wang et al. (2025a) use token accuracy ratio to compare multi-agent collaboration strategies. In contrast, we analyze computational cost against task accuracy to measure Pareto optimality. Based on this, we design ProST, which gives better performance accuracy with higher computational efficiency (more Pareto optimal).
@@ 3 Multi-Agent Design for AppWorld
Multi-agent systems using SLMs instead of LLMs have emerged as an alternative for better cost-performance trade-off Erdogan et al. (2025); Belcak et al. (2025). However, optimizing SLMs for long and complex reasoning tasks remains challenging(Shen et al., 2024). In this work, we focus on AppWorld  (Trivedi et al., 2024), an agent environment of complex tasks requiring coding and ability to interact with APIs. Our method is based on two considerations: (i) SLMs are less effective at essential agent skills like coding, long-context and recovering from errors; (ii) overall computation cost is based on both solution length and low per-token cost of SLMs.
@@ 3.1 Multi-agent Architecture
Solving AppWorld tasks require identifying sub-tasks that can solve each complex problem via writing code and self-reflection Madaan et al. (2023). Agents in multi-agent systems work in collaboration to solve such tasks, where each has its specific responsibilities. Such architectures have been used in Qiao et al. (2024) (plan, tool, and reflect agents) and Shi et al. (2024) (grounding, execution, and review agents). Accordingly, our multi-agent system also comprises of similar elements: an Orchestrator that generates and delegates sub-tasks, an Executor that writes code to interact with AppWorld, and a Critic that provides feedback to Executor.
@@
This agent decomposes the task into subtasks with specific goals and execution plan. Orchestrator operates iteratively, executing one subtask after another and handling execution failures by dynamic decomposition strategy,i.e., further decomposing the subtask (Prasad et al., 2024).
@@
Agent responsible for solving each subtask generated by Orchestrator. It does so by writing code according to the plan, executing in AppWorld and observing the results. Each subtask may need multiple iterations of such cycle. In case of error, Executor refines code. If error still persists it revises the code with feedback from Critic. Execution stops when either task is complete or maximum number of steps is reached.
@@ 4 Progressive Subtask Training
We fine-tune SLMs using multi-agent training trajectories derived from LLMs (described in subsection 5.2). Standard fine-tuning trains each agent to maximize likelihood of the role relevant gold trajectory. However, as we metioned, the trajectories remain long, even with role specialization. SLMs’ inability to proces long trajectories Chen et al. (2025d) limits training them for their respective roles. In particular, we observe that standard finetuning leads to higher error rates for the more difficult middle subtasks compared to first and last subtasks. This observation motivates our proposed solution, drawing from curriculum learning.
@@
This increment, however, may not be equal at each iteration as number of subtasks may not be perfectly distributable across the iterations. Hence, our algorithm tries to optimize subtask increment at later epochs. This is higlighted in Figure 2 where first two epochs have no increment in subtasks but epochs $2$ to $4$ do, as $6$ subtasks cannot be divide into $5$ epochs for equal increment across epochs. (Further details about how subtasks are expanded progressively in Appendix B.3). Below, we formally describe this progressive subtask learning strategy.
@@
In this learning paradigm, we train agents using both correct steps and self-corrected steps. This helps develop a comprehensive understanding of both successful subtask completion and error correction patterns. Following Chen et al. (2025c) we compute the loss function $J^{(e)}_{E}$ (eq. 4), $J^{(e)}_{C}$ (eq. 5) only on correct and self-refined steps for Executor and Critic, respectively. In equation 3, $N_{x}$ is the total number of turns in a given task $x$.
@@ 5.2 Data Synthesis
AppWorld only provides training pairs (input prompts, output state). Hence, to obtain training trajectories, we design a data synthesis pipeline that first generates ReAct Yao et al. (2023)-style trajectories for single agents and then uses a powerful LLM to translate them to multi-agent trajectories. First, we use Llama-3.3-70B-Instruct Meta (2025) to generate single-agent trajectories (see the prompt in Figure 8 in Appendix D) from AppWorld train and dev sets via rejection sampling at different temperatures Chen et al. (2025c), obtaining $2,844$ gold trajectories. Then, for each agent (Orchestrator,Executor and Critic) in our multi-agent setup, we prompt Claude 3.7-Sonnet Anthropic (2025) to convert the single agent trajectories into appropriate trajectories for each agent (See the prompt in Figure 12 in Appendix D). These trajectories were then validated using AppWorld. We gathered $2,708$ such trajectories for training (see Appendix A.1 and A.2 for details).
@@ 5.3 Models & Setup
We use Qwen2.5-Coder variants (7B,14B, and 32B) for our experiments due to their strong coding and reasoning abilities Hui et al. (2024), skills essential for agents. All models are finetuned using LoRA Hu et al. (2021) (see Appendix B.1 for more training and evaluation details). We use the same ReAct Yao et al. (2023) prompt from Trivedi et al. (2024) during inference, but we customize it to be agent-specific (see Appendix D). For each variant, we evaluate five different baselines: single-agent (SA), and fine-tuned version ([FT]), our multi-agent (MA) system proposed in Section 3 and its fine-tuned version ([FT]) and multi-agent system trained with Progressive Subtask Training ([ProST]). We denote agents in multi-agent configurations using the (Orchestrator-Executor-Critic) format, where each placeholder indicates the agent size in billions of parameters (e.g., 7-7-7). We also evaluate the multi-agent system with different sized Orchestrator models, such as a 14B Orchestrator with 7B Executor and Critic (14-7-7), and a 7B Orchestrator with 14B Executor and Critic (7-14-14).
@@ 5.4 Cost-Performance Pareto Analysis
We use Floating Point Operations (FLOPs) to measure computational cost [^3] As we discuss in the limitations section Limitations, we only focus on computational cost and not on memory considerations.. We approximate FLOPsper instance as a function of the model parameters and tokens involved in the computation. We use the formula from Kaplan et al. (2020) which is $2\times\#\text{params}\times(\text{input tokens}+\text{output tokens})$.
Our evaluations plot these effectiveness metrics against FLOPs to show the Pareto curves Yue et al. (2025); Pimentel et al. (2020) of different agent setups, thereby providing Pareto-optimal choices at different cost-performance trade-offs.
@@ 6.1 Effectiveness
. Specializing SLMs for specific roles in a multi-agent setup is expected to improve their effectiveness. Even with no fine-tuning, multi-agent setups are better than single agents. For each parameter variant, the multi-agent system (MA) is better than the corresponding single-agent baseline (SA).
@@
Multi-agent systems finetuned using Progressive Subtask Training ([ProST]) are most effective. Compared to standard fine-tuning, ProST yields a $18.8\%$ relative gain in  for 7B and a $18.5\%$ relative gain for 14B. However, SGC values are the same for ProST and standard finetuning, showing both methods have similar consistency in task solving.
@@ 6.2 Pareto Front Comparisons
Figure 3 shows the computational efficiency (measured in FLOPs) against task accuracy (TGC) of a subset of our systems. It shows three Pareto fronts for non-finetuned, finetuned, and ProST fine-tuned models. The figure supports three key findings:
@@
Recent work shows that powerful planning/Orchestrator agents can improve task effectiveness (Erdogan et al., 2025). We study whether allocating higher capacity to the Orchestrator leads to more Pareto-optimal solutions with ProST (see Figure 4.) To this end, we create two system variants: (i) 7-14-14, which uses a weaker Orchestrator (7B), and (ii) 14-7-7, which uses a stronger Orchestrator (14B).

### v2

@@ Abstract
Biddut Sarker Bijoy1   Mohammad Saqib Hasan1   Pegah Alipoormolabashi1 Avirup Sil2 thanks: Work done while at IBM Research AI.Aruna Balasubramanian1Niranjan Balasubramanian1 1Stony Brook University  2Oracle Applied Science 1{bbijoy, mdshasan, palipoormola, arunab, niranjan}@cs.stonybrook.edu,   2avi.sil@oracle.com
@@ 1 Introduction
Solving complex problems requires a wide-range of capabilities including planning sub-tasks, acting in an environment via coding, reasoning about the actions, recovering from errors, as well as tracking and managing the execution of sub-tasks. LLMs with their strong coding and reasoning abilities and advances in handling long contexts have shown impressive breakthroughs on such complex problems (e.g., AppWorld (appworld-acl24) and SWE-Bench (jimenez2024swebench)). However, these large models incur high computational and API costs for both training and inference. Another viable alternative is to use multi-agent solutions with smaller language models [^1] Small and large are subjective terms. For this paper we use SLM for models with fewer than 40B parameters. (SLMs). The core idea is to identify specialized roles in solving these complex tasks (e.g., orchestration, code writing, and critiquing) and use separate SLMs (or SLM calls) for each role. belcak2025smalllanguagemodelsfuture argue that SLMs are much more cost-effective and practical for training, adapting, and deploying multiple specialized experts for different agentic routines.  shen2024small show these multi-agent solutions with SLMs can be effective for complex problems requiring tool use via APIs.
We consider two challenges that arise in this context. First, for SLMs, even specialized roles can be difficult to learn. Complex problems often have multiple subtasks resulting in long trajectories. Despite their improving capabilities, SLMs still struggle to learn all subtasks under standard fine-tuning. Second, while the per-token compute depends on the size of the models, the overall computational cost also depends upon the nature of their solutions (e.g., their length). Therefore, it is important to explicitly consider both the effectiveness and efficiency of these solutions to better understand the trade-offs.
@@
We evaluate these ideas using Qwen-2.5 Coder, Llama-3.1 and Phi-4 models. We distill training trajectories for a single agent from a frontier model (Claude-3.7-Sonnet). Then, we translate these into role-specific multi-agent training trajectories for three agents, namely, an Orchestrator (planner), an Executor (coding) agent, and a Critic agent. Our evaluations yield multiple interesting observations AppWorld tasks: (i) While larger models tend to have higher effectiveness, multi-agent solutions can achieve better efficiency versus effectiveness trade-offs. (ii) Progressive Subtask Training yields a better Pareto-front compared to standard fine-tuning. (iii) Stronger (higher compute) agents for planning are more effective (and efficient) than using stronger coding and critic agents when finetuning using Progressive Subtask Training. [^2] Code and data available at https://github.com/StonyBrookNLP/prost-multiagents
@@ 2 Related Work
SLMs are being used as agents due to lower computational cost, but struggle with complex planning and tool use (shen2024small), skills essential for agents. Hence, many strategies are proposed: separation of roles to only tackle small, concrete tasks (shi-etal-2024-learning; qiao-etal-2024-autoact); distilling training data from powerful LLMs(shi-etal-2024-learning), Direct Policy Optimization (DPO)(feng2025empoweringllmstaskorienteddialogues); online learning by generating data from unsuccessful attempts(qi2025webrltrainingllmweb), masking loss from erroneous steps(fu2025agentrefineenhancingagentgeneralization); learning only from critical steps (chen2025atlasagenttuninglearning). Contrary to these works, we formulate ProST which optimizes SLMs, as part of multi-agent design, by initiating with important parts of the trajectory and iteratively increasing steps, a strategy never tried for agents.
@@
Many works propose techniques to make multi-agents more robust. These techniques include dynamically reducing redundant agents and messages (wang2025agentdropoutdynamicagentelimination; zhang2024cutcrapeconomicalcommunication), iterative answer refinement through feedback(chen2025magicore), optimizing collaboration strategies(wang-etal-2025-beyond; zeng-etal-2025-s2) and communication protocols (xiao2025longcontextscalingdivide), and test-time search for effective tool planning (chen-etal-2025-smurfs). Recent works propose using cost-effective SLMs for multi-agent setups: erdogan2025planandactimprovingplanningagents uses smaller32-B models as planning and execution subagents; belcak2025smalllanguagemodelsfuture argue for SLM multi-agents based on their improved capabilities, cost, ease of adaptability, and deployment. Our work introduces ProST, a training algorithm for SLMs in multi-agent setups that iteratively trains on subtasks. Unlike previous works, we show how ProST produce SLMs effective at solving tasks as well as computationally efficient.
@@
Recent works have sought to assess and improve the computational efficiency of multi-agent systems. For example, wang2025agentdropoutdynamicagentelimination use token consumption as a metric to prune redundancies in multi-agent collaborations. yue-etal-2025-masrouter show, via Pareto fronts, that their language model routing strategy in multi-agent systems reduces dollar cost while improving accuracy. wang-etal-2025-beyond use token accuracy ratio to compare multi-agent collaboration strategies. In contrast, we analyze computational cost against task accuracy to measure Pareto optimality. Based on this, we design ProST, which gives better performance accuracy with higher computational efficiency (more Pareto optimal).
@@ 3 Multi-Agent Design for AppWorld
Multi-agent systems using SLMs instead of LLMs have emerged as an alternative for better cost-performance trade-off erdogan2025planandactimprovingplanningagents; belcak2025smalllanguagemodelsfuture. However, optimizing SLMs for long and complex reasoning tasks remains challenging(shen2024small). In this work, we focus on AppWorld  (appworld-acl24), an agent environment of complex tasks requiring coding and ability to interact with APIs. Our method is based on two considerations: (i) SLMs are less effective at essential agent skills like coding, long-context and recovering from errors; (ii) overall computation cost is based on both solution length and low per-token cost of SLMs.
@@ 3.1 Multi-agent Architecture
Solving AppWorld tasks require identifying sub-tasks that can solve each complex problem via writing code and self-reflection madaan2023self. Agents in multi-agent systems work in collaboration to solve such tasks, where each has its specific responsibilities. Such architectures have been used in qiao-etal-2024-autoact (plan, tool, and reflect agents) and shi-etal-2024-learning (grounding, execution, and review agents). Accordingly, our multi-agent system (refer to Figure 1) also comprises of similar elements: an Orchestrator that generates and delegates sub-tasks, an Executor that writes code to interact with AppWorld, and a Critic that provides feedback to Executor (see prompt in Figure 10).
@@
This agent decomposes the task into subtasks with specific goals and execution plan. Orchestrator operates iteratively, executing one subtask after another and handling execution failures by dynamic decomposition strategy,i.e., further decomposing the subtask (prasad2024adaptasneededdecompositionplanning).
@@
Agent responsible for solving each subtask generated by Orchestrator. It does so by writing code according to the plan, executing in AppWorld and observing the results (see prompt in Figure 11). Each subtask may need multiple iterations of such cycle. In case of error, Executor refines code. If error still persists it revises the code with feedback from Critic. Execution stops when either task is complete or maximum number of steps is reached.
@@ 4 Progressive Subtask Training
We fine-tune SLMs using multi-agent training trajectories derived from LLMs (described in subsection 5.2). Standard fine-tuning trains each agent to maximize likelihood of the role relevant gold trajectory. However, as we metioned, the trajectories remain long, even with role specialization. SLMs’ inability to process long trajectories chen2025atlasagenttuninglearning limits training them for their respective roles. In particular, we observe that standard finetuning leads to higher error rates for the more difficult middle subtasks compared to first and last subtasks. This observation motivates our proposed solution, drawing from curriculum learning.
@@
This increment, however, may not be equal at each iteration as number of subtasks may not be perfectly distributable across the iterations. Hence, our algorithm tries to optimize subtask increment at later epochs. This is highlighted in Figure 2 where first two epochs have no increment in subtasks but epochs $2$ to $4$ do, as $6$ subtasks cannot be divide into $5$ epochs for equal increment across epochs. (Further details about how subtasks are expanded progressively in Appendix B.3). Below, we formally describe this progressive subtask learning strategy.
@@
In this learning paradigm, we train agents using both correct steps and self-corrected steps. This helps develop a comprehensive understanding of both successful subtask completion and error correction patterns. Following fu2025agentrefineenhancingagentgeneralization we compute the loss function $J^{(e)}_{E}$ (eq. 4), $J^{(e)}_{C}$ (eq. 5) only on correct and self-refined steps for Executor and Critic, respectively. In equation 3, $N_{x}$ is the total number of turns in a given task $x$.
@@ 5.2 Data Synthesis
AppWorld only provides training pairs (input prompts, output state). Hence, to obtain training trajectories, we design a data synthesis pipeline similar to (erdogan2025planandactimprovingplanningagents) that first generates ReAct react-style trajectories for single agents and then uses a powerful LLM to translate them to multi-agent trajectories. First, we use Llama-3.3-70B-Instruct llama2025instruct33 to generate single-agent trajectories (see the prompt in Figure 13 in Appendix E) from AppWorld train and dev sets via rejection sampling at different temperatures chen2025reinforcementlearninglonghorizoninteractive, obtaining $2,844$ gold trajectories. Then, for each agent (Orchestrator,Executor and Critic) in our multi-agent setup, we prompt Claude 3.7-Sonnet anthropic2025claude37 to convert the single agent trajectories into appropriate trajectories for each agent (See the prompt in Figure 14 in Appendix E). These trajectories were then validated using AppWorld. We gathered $2,708$ such trajectories for training (see Appendix A.1 and A.2 for details).
@@ 5.3 Models & Setup
We use Qwen2.5-Coder variants (7B,14B, and 32B) for our experiments due to their strong coding abilities and Llama-3.1-8B and Phi-4 for their strong reasoning abilities hui2024qwen25codertechnicalreport, skills essential for agents. All models are finetuned using LoRA hu2021loralowrankadaptationlarge (see Appendix B.1 for more training and evaluation details). We use the same ReAct react prompt from appworld-acl24 during inference, but we customize it to be agent-specific (see Appendix E). For each variant, we evaluate five different baselines: single-agent (SA), and fine-tuned version ([FT]), our multi-agent (MA) system proposed in Section 3 and its fine-tuned version ([FT]) and multi-agent system trained with Progressive Subtask Training ([ProST]). We denote agents in multi-agent configurations using the (Orchestrator-Executor-Critic) format, where each placeholder indicates the agent size in billions of parameters (e.g., 7-7-7). We also evaluate the multi-agent system with different sized Orchestrator models, such as a 14B Orchestrator with 7B Executor and Critic (14-7-7), and a 7B Orchestrator with 14B Executor and Critic (7-14-14).
@@ 5.4 Cost-Performance Pareto Analysis
We use Floating Point Operations (FLOPs) to measure computational cost [^4] As we discuss in the limitations section Limitations, we only focus on computational cost and not on memory considerations.. We approximate FLOPsper instance as a function of the model parameters and tokens involved in the computation. We use the formula from kaplan2020scaling which is $2\times\#\text{params}\times(\text{input tokens}+\text{output tokens})$.
Our evaluations plot these effectiveness metrics against FLOPs to show the Pareto curves yue-etal-2025-masrouter; pimentel-etal-2020-pareto of different agent setups, thereby providing Pareto-optimal choices at different cost-performance trade-offs.
@@ 6.1 Effectiveness
We expect multi-agent systems with specialized SLMs than their corresponding single agent versions. This is observed in Table 1 and 3 as the multi-agent system (MA) is better than the corresponding single-agent baseline (SA) for each parameter variant (e.g., MA 14-14-14 [ProST] attains TGC score of $42.3$, while SA 14 [FT] TGC scores $34.5$). Furthermore, we find that multi-agent systems can even be on par with large single agents (e.g., GPT-4o) in terms of both effectiveness as well as efficiency [^5] Here we mean efficiency in terms of overall FLOPs and not accounting of engineering overheads that can impact both settings.. For example, the Phi-4 14B-14B-14B multi-agent gets 46.4, which is close to GPT-4o gets 48.8 on Test-Normal TGC) appworld-acl24.
@@
To solve AppWorld tasks, we need models that are good at both complex reasoning as well as coding. Multi-agents powered by code specialist SLMs as base models (e.g., Qwen-2.5-Coder) improve when trained with ProST. However, as shown in Table 1 and 3, when trained with models that are good at both coding and complex reasoning, such as Llama-3.1 and Phi-4 base models, we find that ProST improves SGC scores on three out of four models, demonstrating its utility for improving robustness as well. Compared to standard fine-tuning, ProST yields $18.8\%$ and $18.5\%$ relative gains in TGC scores for the 7B and 14B coding models, respectively. Using Llama-3.1 and Phi-4, the relative gains are even larger at $30.8\%$ and $34.5\%$, respectively.
@@ 6.2 Pareto Front Comparisons
Figure 3 shows the computational efficiency (measured in FLOPs) against task accuracy (TGC) of a subset of our systems using Qwen-2.5-Coder series models. It shows three Pareto fronts for non-finetuned, finetuned, and ProST fine-tuned models. The figure supports three key findings:
@@
Recent work shows that powerful planning/Orchestrator agents can improve task effectiveness (erdogan2025planandactimprovingplanningagents). We study whether allocating higher capacity to the Orchestrator leads to more Pareto-optimal solutions with ProST (see Figure 4.) To this end, we create two system variants: (i) 7-14-14, which uses a weaker Orchestrator (7B), and (ii) 14-7-7, which uses a stronger Orchestrator (14B).